OpenStudy (anonymous):

For each sum find: a.) the number of terms, b.) the first term, c.) the last term, and d.) evaluate the sum.

OpenStudy (anonymous):

\[\sum_{n=1}^{5} (3n+1)\] \[\sum_{n=1}^{8} 2n/3\]

OpenStudy (stamp):

a: 5 terms, 8 terms b: 3(1)+1, 2(1)/3 c: 3(5)+1, 2(8)/3 d:

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

idk how!!!!!!!!!!!!!!!!!!!

OpenStudy (stamp):

\[n_1=4,\ n_2=7,\ n_3=10,\ ...\]This is an arithmetic series. The sum of an arithmetic series is described by the equation\[\frac{n}{2}(2a+(n-1)d)\]n is the number of terms, a is the first term, and d is the difference between each term. In our case, n =5, a = 4, and d = 3.\[\frac{5}{2}(2(4)+(5-1)3)=50\] You can check. 4 + 7 + 10 + 13 + 16 = 50. So the sum is 50.